OpenStudy (anonymous):

What is the real part of (cos30 +isin30)^10

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[(cos(30)+isin(30))^{10}=cos(300)+isin(300)\]

OpenStudy (anonymous):

\[= \frac{1}{2}-\frac{\sqrt{3}}{2}i\]

OpenStudy (anonymous):

real part is \[\frac{1}{2}\] imaginary part is \[-\frac{\sqrt{3}}{2}\]